OpenStudy (anonymous):

I posted this question like 5 times! Please explain to me how to get the answer! :( Luke shaped a piece of wire that was 6/11 meter long to form a circle. He divided the length of the wire by 22/7 to find that the diameter of the circle was 7/12 meter. Which of the following statements explains the error Luke made while calculating the diameter of the circle? A. He used the reciprocal of 6/11 while multiplying the two fractions. B. He did not find the reciprocal of 22/7 while multiplying the two fractions. C. He used the reciprocals of 6/11 and 22/7 while multiplying the two fractions

OpenStudy (anonymous):

D. He did not find the reciprocal of the answer after multiplying the two fractions.

OpenStudy (anonymous):

Test each case individually

OpenStudy (amistre64):

C = d pi for starters

OpenStudy (amistre64):

C/pi = d

OpenStudy (amistre64):

22/7 is about pi

OpenStudy (anonymous):

case a: (11/6)*(7/22)=7/12 looks like we can stop here. This is where he messed up

OpenStudy (amistre64):

6/11 ---- = 7/12 will give you what he did i beleive x

OpenStudy (anonymous):

Thank you! Finally I get some help...

OpenStudy (amistre64):

x = 72/77 in that case .. which doesnt seem to be an answer x ----- = 7/12 then like dw said 22/7

OpenStudy (amistre64):

x = 11/6 in that case so from the options a does seem most probable

OpenStudy (anonymous):

Actually, i think the answer is option c because in a it implies he multiplied by pi instead of the reciprocal

OpenStudy (amistre64):

its hard to tell what he did ....

OpenStudy (amistre64):

you dont multiply fractions ...

OpenStudy (anonymous):

Okay, Dockworker. I was going to go with Choice C anyway...

OpenStudy (amistre64):

you do, but not in this case so its a bit ackward to read the options

OpenStudy (anonymous):

yes...the options are kind of vague and then a makes no sense when you get to c

OpenStudy (amistre64):

c is best tho

OpenStudy (amistre64):

he flipped both fractions and multiplied instead of jsut flipping the bottom one

OpenStudy (anonymous):

11/6 ÷ 22/7 which gets you 7/12.

OpenStudy (anonymous):

i assume when they said multiplying fractions that he was doing it properly in case a, but the question is dumb

OpenStudy (amistre64):

its spose to be designed to get you to consider how people might mess up; but yeah ... its a rather poor question

OpenStudy (anonymous):

I'll just go with choice C. Thanks for the help!